OpenStudy (anonymous):

n a song competition, a participant has to score a total of at least 20 points in the first four rounds combined to move on to the fifth and final round. Curtis scored 2 points in the first round. He then went on to score additional points in the second, third, and fourth rounds. In each of those rounds, his score was identical. Which inequality best shows the number of points, p, that Curtis scored in each of the second, third, and fourth rounds if he earned a place in the finals?

OpenStudy (anonymous):

@AaronAndyson

OpenStudy (anonymous):

do you understand this

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

k

OpenStudy (anonymous):

do you get this

OpenStudy (danjs):

\[Total ~ Points \ge 20\]

OpenStudy (danjs):

First Round - 2 2,3,4 Rounds - p \[2 + 3p \ge 20\] to make it to fifth round

OpenStudy (anonymous):

so that would be the equation

OpenStudy (danjs):

yeah, well, solve for p

OpenStudy (anonymous):

hold on im solving it right now

OpenStudy (anonymous):

what should i get

OpenStudy (danjs):

...

OpenStudy (danjs):

p is greater than or equal to six

OpenStudy (danjs):

makes sense, to just make fifth round he has to get 20... 2 + 6 + 6 + 6 = 20

OpenStudy (anonymous):

i get it now